# Parametric pair production of collective excitations in a Bose-Einstein condensate

### v1

@@ Acknowledgments
This article was submitted to the special issue of Comptes Rendus Physique honoring Renaud Parentani. We warmly dedicate this work to him and gratefully acknowledge his role in inspiring and initiating this research. We acknowledge funding from QuantERA Grant No. ANR-22-QUA2-000801 (MENTA) and ANR Grant No. 20-CE-47-0001-01 (COSQUA), the LabEx PALM (ANR-10-LABX-0039PALM). V.G. was funded by the Région Ile-de-France in the framework of the DIM SIRTEQ program and the Quantum Saclay program FQPS (ANR-21-CMAQ-0002). R.D. acknowledges a Ph.D grant with reference 2024.03181.BD from the Portuguese Foundation for Science and Technology (FCT). S.R. is funded by the CNRS Chair in Physical Hydrodynamics (ANR-22-CPJ2-0039-01).
@@ Appendix A Model
Neglecting transverse excitations, we integrate over this profile to have an effective one-dimensional gas with a contact interaction $g_{1}=2\hbar a_{s}^{2}/(m\sigma^{2})$  [^4] Even without assuming a Gaussian profile for the transverse density, a similar result can be obtained [37].. The small excitations around the BEC of our weakly interacting gas are well described within Bogoliubov theory [78]. We split the atomic field $\hat{\Psi}(z)$ into a classical piece $\Psi_{0}$ describing the homogeneous BEC, and a quantum perturbation $\delta\hat{\Psi}(z)$ describing the uncondensed atoms, and truncate the Hamiltonian at second order in this perturbation $H^{(2)}$ [79]  [^5] The difficulties with Bogoliubov treatment for a 1D homogenous gas [81] are irrelevant here since we focus on a pair of modes only.. The dynamics of these interacting uncondensed atoms can be conveniently recast as that of freely evolving collective excitations, or quasiparticles. The are represented by annihilation (creation) operators $\hat{b}_{k}^{(\dagger)}$ related to the atomic one $\hat{a}_{k}^{(\dagger)}$ by a Bogoliubov transformation $\hat{a}_{k}=u_{k}\hat{b}_{k}+v_{k}\hat{b}_{-k}^{\dagger}$, where the values of $u_{k},v_{k}$ are given in the main text. The quasiparticle operators by definition diagonalize the Hamiltonian $H^{(2)}=\sum_{k\neq 0}\hbar\omega_{k}(g_{1})\hat{b}_{k}^{\dagger}\hat{b}_{k}$ with $\omega_{k}$ given by Eq. (1) and $n_{1}=\lvert\Psi_{0}\rvert^{2}$ the density of the BEC. Thus, a collective excitation of momentum $k$ has an energy $\hbar\omega_{k}$ and consists of atoms with opposite momenta $\pm k$ that keep interacting. Our experimental set-up allows us to measure atom numbers $n^{\rm at}_{k}=\langle\hat{a}_{k}^{\dagger}\hat{a}_{k}\rangle$ as a function of momentum $k$ which is related to the number of quasiparticles $n_{k}=\langle\hat{b}_{k}^{\dagger}\hat{b}_{k}\rangle$ and their anomalous correlations $c_{k}=\langle\hat{b}_{k}\hat{b}_{-k}\rangle$ by Eq. (5).
@@
where $\alpha\left(t;t^{\prime}\right),\beta\left(t;t^{\prime}\right)$ are solutions of Eq. (1) with initial conditions $\alpha\left(t^{\prime};t^{\prime}\right)=1,\beta\left(t^{\prime};t^{\prime}\right)=0$. At the start of the experiment the trap is held constant $\omega_{\perp}=\omega_{\perp,0}$, the gas’s transverse size $\sigma$ is fixed and so $\omega_{k}$ is time-independent, which leaves only the first term in Eq. (1) encoding the free evolution $\hat{b}_{k}(t)=\hat{b}_{k}(0)e^{-i\omega_{k}t}$. The quasiparticle density is time-independent $n_{k}(t)=n_{k}(0)$ and the correlation oscillates $c_{k}(t)=e^{-i2\omega_{k}t}c_{k}(0)$ at $2\omega_{k}$. When the trap frequency $\omega_{\perp}$ is varied, the transverse size of the gas reacts according to Eq. (7) and $\omega_{k}$ becomes time-dependent. The second term in Eq.(1) then does lead to a production of pairs of opposite momentum quasiparticles which can be separated in the following two stages [37].
@@ A.1 Modulation
We initially modulate the trap at resonance $\omega_{\perp,0}$ for $\delta t=4\times\pi/\omega_{\perp,0}$ after which we hold the trap at $\omega_{\perp,0}$ for a duration $\Delta t$. Eq. (7) then predicts that $\sigma$ first experiences breathing oscillations at frequency $2\omega_{\perp,0}$ with linearly growing amplitude, which then persist at fixed amplitude during the hold time, see Fig. 1. This variation leads $g_{1}$ to oscillate at $2\omega_{\perp,0}$ [37]. To get explicit expressions, we neglect the growth phase and assume sinusoidal oscillation with a fixed amplitude $a$ i.e. $g_{1}\left(t\right)$ is given by Eq. (3) from $t=0$ to $\Delta t$. This modulation corresponds to a two-mode squeezing operation on the $\pm k$ modes which are resonant with the process $\omega_{\pm k}=\omega_{\perp,0}$ [36]. Assuming an initial thermal population $n_{\pm k}(0)=n_{k}^{\mathrm{in}}$ and $c_{k}(0)=0$, Eq. (8) can be solved using a rotating wave approximation and gives [36]
@@
[…]
[width=]fig5.pdf
FIG. 5: a, Quasiparticle and atomic populations during the excitation. b, Corresponding evolution of the real part of the anomalous correlation $\mathrm{Re}(c_{k})$ (left $y$ scale, solid yellow) and of the Bogoliubov coefficients $u_{k}$ and $v_{k}$ (right $y$ scale, in blue dashed).
@@ A.3 Analytical growth of the atom number
We assume that $g_{1}$ is modulated sinusoidally as Eq. (3) for a duration $\Delta t$ before opening the trap. The number and correlation of quasiparticles are then given by Eq. (9) at opening time. The detected number of atoms $n_{k}^{\rm det}$ in the resonant mode $\omega_{k}=\omega_{d}/2$ after trap expansion is obtained by inserting these function in Eq. (11) and multiplying the overall number atom number by $\eta$ the quantum efficiency of the detector. We then obtain
@@ A.4 Numerical simulation of both phenomena
The predicted amplitude in Eq. (14) relies on the value of $\lvert\beta_{\rm op}\rvert$, which characterizes the Bogoliubov transformation from the quasiparticle to the atomic basis. We can then simulate numerically the expected amplitude within our model. Using our Gaussian ansatz, we infer $g_{1}n_{1}$ and using the dispersion relation we find the resonant modes $k$ such that $\omega_{k}=\omega_{\perp}$. We then numerically solve the full dynamics of $\sigma$ given by Eq. (7): an initial four periods of modulation at $2\omega_{\perp,0}$, a hold time of $\Delta t$, and then the opening. Using this variation we numerically solve Eq. (8) from an initial thermal state of quasiparticles at temperature $T=35$ nK, to a late enough time after the trap switch-off such that the Bogoliubov coefficients have become time-independent. This gives us access numerically to the number of detected atoms Eq. (2), with a 25% quantum efficiency. We repeat this procedure for two different switch-off ramps, a very short one of $0.1\,\mathrm{ms}$ and a longer one of $1\,\mathrm{ms}$, matching that of the experimental realizations shown in Fig. 4b. The corresponding evolution of the atom number is shown in Fig. 4a. The solid red line corresponds to the case where the dipole trap power is switched off abruptly and the dashed green curve assumes that the dipole trap power is ramped down in 1.5 ms. In agreement with the experimental data, we observe pronounced oscillations of relative amplitude $A_{k}\approx 0.34$ for the shortest ramp, roughly half that of the in situ oscillations for which Eq. (6) gives $A_{k}^{\mathrm{inst}}\approx 0.7$, and much reduced oscillation for the slowest turn off, with an amplitude of 0.1. As for the experimental data, we observe that the green dashed curve is above the solid red one, an indication of the additional quasiparticle production that occurs during the slow switch-off.

### v2

@@ Acknowledgments
We thank and dedicate this work to Renaud Parentani, to whom we owe the genesis of this study. We acknowledge funding from QuantERA Grant No. ANR-22-QUA2-000801 (MENTA) and ANR Grant No. 20-CE-47-0001-01 (COSQUA), the LabEx PALM (ANR-10-LABX-0039PALM). V.G. was funded by the Région Ile-de-France in the framework of the DIM SIRTEQ program and the Quantum Saclay program FQPS (ANR-21-CMAQ-0002). R.D. acknowledges a Ph.D grant with reference 2024.03181.BD from the Portuguese Foundation for Science and Technology (FCT). S.R. is funded by the CNRS Chair in Physical Hydrodynamics (ANR-22-CPJ2-0039-01).
@@ Appendix A Model
Neglecting transverse excitations, we integrate over this profile to have an effective one-dimensional gas with a contact interaction $g_{1}=2\hbar a_{s}^{2}/(m\sigma^{2})$  [^4] Even without assuming a Gaussian profile for the transverse density, a similar result can be obtained [37].. The small excitations around the BEC of our weakly interacting gas are well described within Bogoliubov theory [78]. We split the atomic field $\hat{\Psi}(z)$ into a classical piece $\Psi_{0}$ describing the homogeneous BEC, and a quantum perturbation $\delta\hat{\Psi}(z)$ describing the uncondensed atoms, and truncate the Hamiltonian at second order in this perturbation $H^{(2)}$ [79]  [^5] The difficulties with Bogoliubov treatment for a 1D homogenous gas [81] are irrelevant here since we focus on a pair of modes only.. The dynamics of these interacting uncondensed atoms can be conveniently recast as that of freely evolving collective excitations, or quasiparticles. They are represented by annihilation (creation) operators $\hat{b}_{k}^{(\dagger)}$ related to the atomic one $\hat{a}_{k}^{(\dagger)}$ by a Bogoliubov transformation $\hat{a}_{k}=u_{k}\hat{b}_{k}+v_{k}\hat{b}_{-k}^{\dagger}$, where the values of $u_{k},v_{k}$ are given in the main text. The quasiparticle operators by definition diagonalize the Hamiltonian $H^{(2)}=\sum_{k\neq 0}\hbar\omega_{k}(g_{1})\hat{b}_{k}^{\dagger}\hat{b}_{k}$ with $\omega_{k}$ given by Eq. (1) and $n_{1}=\lvert\Psi_{0}\rvert^{2}$ the density of the BEC. Thus, a collective excitation of momentum $k$ has an energy $\hbar\omega_{k}$ and consists of atoms with opposite momenta $\pm k$ that keep interacting. Our experimental set-up allows us to measure atom numbers $n^{\rm at}_{k}=\langle\hat{a}_{k}^{\dagger}\hat{a}_{k}\rangle$ as a function of momentum $k$ which is related to the number of quasiparticles $n_{k}=\langle\hat{b}_{k}^{\dagger}\hat{b}_{k}\rangle$ and their anomalous correlations $c_{k}=\langle\hat{b}_{k}\hat{b}_{-k}\rangle$ by Eq. (5).
@@
where $\alpha\left(t;t^{\prime}\right),\beta\left(t;t^{\prime}\right)$ are solutions of Eq. (1) with initial conditions $\alpha\left(t^{\prime};t^{\prime}\right)=1,\beta\left(t^{\prime};t^{\prime}\right)=0$. At the start of the experiment the trap is held constant $\omega_{\perp}=\omega_{\perp,0}$, the gas’s transverse size $\sigma$ is fixed and so $\omega_{k}$ is time-independent, which leaves only the first term in Eq. (1) encoding the free evolution $\hat{b}_{k}(t)=\hat{b}_{k}(0)e^{-i\omega_{k}t}$. The quasiparticle number is time-independent $n_{k}(t)=n_{k}(0)$ and the correlation oscillates $c_{k}(t)=e^{-i2\omega_{k}t}c_{k}(0)$ at $2\omega_{k}$. When the trap frequency $\omega_{\perp}$ is varied, the transverse size of the gas reacts according to Eq. (7) and $\omega_{k}$ becomes time-dependent. The second term in Eq.(1) then does lead to a production of pairs of opposite momentum quasiparticles which can be separated in the following two stages [37].
@@ A.1 Modulation
We initially modulate the trap frequency at resonance $2\omega_{\perp,0}$ for $\delta t=4\times\pi/\omega_{\perp,0}$ after which we hold the trap at $\omega_{\perp,0}$ for a duration $\Delta t$. Eq. (7) then predicts that $\sigma$ first experiences breathing oscillations at frequency $2\omega_{\perp,0}$ with linearly growing amplitude, which then persist at fixed amplitude during the hold time, see Fig. 1. This variation leads $g_{1}$ to oscillate at $2\omega_{\perp,0}$ [37]. To get explicit expressions, we neglect the growth phase and assume sinusoidal oscillation with a fixed amplitude $a$ i.e. $g_{1}\left(t\right)$ is given by Eq. (3) from $t=0$ to $\Delta t$. This modulation corresponds to a two-mode squeezing operation on the $\pm k$ modes which are resonant with the process $\omega_{\pm k}=\omega_{\perp,0}$ [36]. Assuming an initial thermal population $n_{\pm k}(0)=n_{k}^{\mathrm{in}}$ and $c_{k}(0)=0$, Eq. (8) can be solved using a rotating wave approximation and gives [36]
@@
Figure 5: Theoretical evolution of the mean value of some operators during the parametric growth. a, Quasiparticle and atomic populations during the excitation. b, Corresponding evolution of the real part of the anomalous correlation $\mathrm{Re}(c_{k})$ (left $y$ scale, solid yellow) and of the Bogoliubov coefficients $u_{k}$ and $v_{k}$ (right $y$ scale, in dashed blue).
@@ A.3 Analytical growth of the atom number
We assume that $g_{1}$ is modulated sinusoidally as Eq. (3) for a duration $\Delta t$ before opening the trap. The number and correlation of quasiparticles are then given by Eq. (9) at opening time. The detected number of atoms $n_{k}^{\rm det}$ in the resonant mode $\omega_{k}=\omega_{d}/2$ after trap expansion is obtained by inserting these function in Eq. (11) and multiplying the overall atom number by $\eta$ the quantum efficiency of the detector. We then obtain
@@ A.4 Numerical simulation of both phenomena
The predicted amplitude in Eq. (14) relies on the value of $\lvert\beta_{\rm op}\rvert$, which characterizes the Bogoliubov transformation from the quasiparticle to the atomic basis. We can then simulate numerically the expected amplitude within our model. Using our Gaussian ansatz, we infer $g_{1}n_{1}$ and using the dispersion relation we find the resonant modes $k$ such that $\omega_{k}=\omega_{\perp}$. We then numerically solve the full dynamics of $\sigma$ given by Eq. (7): an initial four periods of modulation at $2\omega_{\perp,0}$, a hold time of $\Delta t$, and then the opening. Using this variation we numerically solve Eq. (8) from an initial thermal state of quasiparticles at temperature $T=35$ nK, to a late enough time after the trap switch-off such that the Bogoliubov coefficients have become time-independent. This gives us access numerically to the number of detected atoms Eq. (2), with a 25% quantum efficiency. We repeat this procedure for two different switch-off ramps, a very short one of 10 µs and a longer one of 1.5 ms, matching that of the experimental realizations shown in Fig. 4b. The corresponding evolution of the atom number is shown in Fig. 4a. The solid red line corresponds to the case where the dipole trap power is switched off abruptly and the dashed green curve assumes that the dipole trap power is ramped down in 1.5 ms. In agreement with the experimental data, we observe pronounced oscillations of relative amplitude $A_{k}\approx 0.34$ for the shortest ramp, roughly half that of the in situ oscillations for which Eq. (6) gives $A_{k}^{\mathrm{inst}}\approx 0.7$, and much reduced oscillation for the slowest turn off, with an amplitude of 0.1. As for the experimental data, we observe that the green dashed curve is above the solid red one, an indication of the additional quasiparticle production that occurs during the slow switch-off.